### v1

@@ 2.2 ViTs
Due to quadratic complexity of attention in ViT [15], efficient token mixing and global attention approximations have been actively studied with various ideas stemming from shifted window of Swin [25]. In lieu of window shifting, competitive performances have been reported on ImageNet-1K and downstream tasks by techniques such as pooling (Poolformer [45]), shuffling (Shuffleformer [14], mixing across windows and dimensions (Mixformer [3]), high/low frequency global attention decomposition (Hiloformer [32]), pale shaped window attention (Paleformer, [41]), cross shaped windown attention (CSwinformer [5]) among others. See also hybrid and unified CNN–ViT models [40, 12, 8, 43, 7, 20, 19].
@@ 2.4 Lightweight Vision Networks
Lightweight networks are desirable for mobile deployment and resource constrained applications. Separable (group) convolutions and shuffle operations are effective techniques for designing CNNs in the lightweight category, see MobileNets [34], ShuffleNets [47, 27, 26] and references therein among others. Lightweight ViTs have been proposed combining MobileNet and efficient attention blocks in [21, 30, 44] and elsewhere. A lightweight Fourier transform based attention-free vision network is AFF [13] which forms the baseline of our work here.
@@ 3.2.2 Alternating Fourier-Image Domain Filtering
By integrating spatial and Fourier filters, we enhance the AFF architecture via a local-global approximation structure, enabling it to effectively and accurately extract significant visual features. This dual approach ensures that we maintain computational efficiency while achieving the desired adaptive filtering outcomes. Figures 2(b) and 2(a) illustrate this concept, showing the block view and its repetitive three-stage process from input to output, respectively.
@@ 4.1 Image Classification
4.1.1 Dataset
ImageNet-1K [36] is a large-scale dataset containing over 1.2 million images across 1000 object categories for image classification. The dataset features images of varying sizes and resolutions, depicting objects in diverse backgrounds. We train our models on ImageNet-1K from scratch to demonstrate the effectiveness and efficiency of our proposed models in image classification.
4.1.2 Setting
AFIDAF network is trained from scratch for 300 epochs using images with a resolution of $256\times 256$ pixels on 8 NVIDIA RTX A6000 GPUs with a batch size of 1024. We use a cosine learning rate scheduler with a base learning rate of $2\mathrm{e}{-3}$ and a minimum learning rate of $2\mathrm{e}{-4}$. The optimizer is AdamW with a weight decay of 0.05, and the loss function is cross-entropy. Model configurations are detailed below on the architecture of AFIDAF and its tiny version AFIDAF-T derived from AFF-T. Input feature from data prepossessing has size of $256^{2}$ with 1 block and 16 output channels. Then it pass through 1 Conv Stem Layer, 3 Down Sample AFIDAF Block (as Fig.2(a)) and output. The Conv Stem Layer has an output size of $64^{2}$, 4 blocks, and 32 output channels. The First Down Sample AFIDAF Block has an output size of $32^{2}$, 2 blocks, with 96 (128) channels for AFIDAF-T (AFIDAF). The Second Down Sample AFIDAF Block features an output size of $16^{2}$, 4 blocks, with 160 (256) channels for AFIDAF-T (AFIDAF). The Third Down Sample AFIDAF Block has an output size of $8^{2}$, 3 blocks, with 192 (320) channels for AFIDAF-T (AFIDAF).
@@ 4.1.3 Results
The hierarchical AFIDAF model, derived from Swin-T and further compressed, significantly reduces the parameters (from 28M to 14.8M) and FLOPs (from 4.5G to 4.45G) off the original Swin-T, while maintaining competitive performance vs. recent ViTs of similar/larger sizes, with a 1.4 (0.5) % decrease in Top-1 (5) accuracy (refer to Tab. 2).
@@
Table 1: Lightweight network classification comparison on ImageNet-1K dataset. IDAF (using image domain adaptive filtering block) replaces AFF block’s channel mixing with image domain LKA [10].
Model Params (M) Flops (G) Top-1 (%) Top-5 (%)
Swin-T[25] 28 4.5 81.2 95.5
SpectFormer-XS[33] 20 4.0 80.2 94.7
PoolFormer-S12[45] 12 1.8 77.2 -
PoolFormer-S24[45] 21 3.4 80.3 -
GFNet-XS[35] 16 2.9 78.6 94.2
HAFIDAF 14.8 4.45 79.8 95.0
Table 2: Middleweight network classification comparison on ImageNet-1K dataset, with HAFIDAT being a compressed Swin-T [25].
4.1.4 Ablation Study
The last 3 lines of Tab. 1 show the advantage of alternating Fourier and image domain filtering over AFFNet [13] and IDAF (i.e. replace AFF block with image domain LKA [10]) on ImageNet-1K.
@@ 4.2 Object Detection
(a) Comparison of AFIDAF variants.
Model Param(M) APbox(%) AP50box(%) AP75box(%)
R-50[11] 82 46.3 64.3 50.5
DeiT-S[37] 80 48.0 67.2 51.7
Swin-T[25] 86 50.5 69.3 54.9
HAFIDAF 72 48.9 67.6 53.4
(b) Comparison of hierarchical AFIDAF with larger parameter size models.
Table 3: Object detection on MS-COCO 2017 dataset.
4.2.1 Dataset
4.2.2 Setting
For model compression, we consider the backbone model Hierarchical AFIDAF equipped with Cascade Mask R-CNN framework on mmdetection3(b). For these four frameworks, we utilize the same settings: multi-scale training [25] (resizing the input such that the shorter side is between 480 and 800 while the longer side is at most 1333), AdamW optimizer (initial learning rate of $1\mathrm{e}{-4}$, weight decay of 0.05, batch size of 16, 3x schedule (36 epochs), and ImageNet-1K pre-trained model as initialization.
4.2.3 Results
As shown in Tab. (3(a)), the detection models equipped with AFIDAF outperform other lightweight transformer-based detectors in terms of mean Average Precision (mAP). Specifically, AFIDAF surpasses the second-best AFFNet [13] by 1.8% in mAP. Consistently, AFIDAF-T edges out AFFNet-T by 0.1% in mAP with 0.1M more parameters. In Tab. (3(b)), HAFIDAF compresses Swin-T by 17% in parameter size while keeping accuracy drop at 1.5% in AP (box).
4.3 Semantic Segmentation
@@
(a) AFIDAF variants vs. other LW backbones.
Model Param (M) mIoU(%) mAcc(%) aAcc(%)
@@
(b) Comparing hierarchical AFIDAF with Swin-T baseline.
Table 4: Semantic segmentation on PASCAL VOC 2012 dataset.
4.3.1 Dataset
4.3.2 Setting
We use the DeepLabv3 [2] framework for semantic segmentation with AFIDAF and AFIDAF-T backbones. Images are resized to $512\times 512$, and models are initialized with ImageNet-1K pretrained weights. Models are trained for 50 epochs on the VOC dataset, using a cosine learning rate scheduler with a base rate of $5\mathrm{e}{-4}$, a minimum rate of $1\mathrm{e}{-6}$, and optimizer AdamW with a weight decay of 0.05. The loss function employed is cross-entropy loss. In training HAFIDAF models, we utilize UperNet in mmseg[4] as our base framework for model compression.
4.3.3 Results
In Tab. (4(a)), AFIDAF demonstrates superior performance compared to other lightweight networks for semantic segmentation. Specifically, AFIDAF achieves a mean Intersection over Union (mIoU) of 81.6%, surpassing the second-best lightweight network, AFFNet, by 1.1%. Additionally, AFIDAF-T exceeds AFFNet-T by 1.8% in mIoU. In Tab. (4(b)), HAFIDAF compresses Swin-T by 24% and improves all three measures of accuracy. These results highlight the effectiveness of our proposed method for dense prediction tasks.

### v2

@@ 2.2 ViTs
Due to quadratic complexity of attention in ViT [14], efficient token mixing and global attention approximations have been actively studied with various ideas stemming from shifted window of Swin [24]. In lieu of window shifting, competitive performances have been reported on ImageNet-1K and downstream tasks by techniques such as pooling (Poolformer [45]), shuffling (Shuffleformer [13]), mixing across windows and dimensions (Mixformer [3]), high/low frequency global attention decomposition (Hiloformer [31]), pale shaped window attention (Paleformer, [41]), cross shaped windown attention (CSwinformer [4]) among others. See also hybrid and unified CNN–ViT models [40, 11, 7, 43, 6, 19, 18].
@@ 2.4 Lightweight Vision Networks
Lightweight networks are desirable for mobile deployment and resource constrained applications. Separable (group) convolutions and shuffle operations are effective techniques for designing CNNs in the lightweight category, see MobileNets [33], ShuffleNets [47, 26, 25] and references therein among others. Lightweight ViTs have been proposed combining MobileNet and efficient attention blocks in [20, 29, 44], see also [38] for a ViT motivated mobile CNN. A lightweight Fourier transform based attention-free vision network is AFF [12] which forms the baseline of our work here.
@@ 3.2.2 Alternating Fourier-Image Domain Filtering
By integrating spatial and Fourier filters, we enhance the AFF architecture via a local-global approximation structure, enabling it to effectively and accurately extract significant visual features. This dual approach ensures that we maintain computational efficiency while achieving the desired adaptive filtering outcomes. Fig. 2(b) and 2(a) illustrate this concept, showing the block view and its repetitive three-stage process from input to output, respectively.
@@ 4.1 Image Classification
Setting
The ImageNet-1K dataset [35], containing over 1.2 million images across 1000 object categories, is utilized for training our models from scratch to validate the effectiveness and efficiency of our proposed AFIDAF network in image classification. We trained AFIDAF from scratch for 300 epochs using $256\times 256$ pixel images on 8 NVIDIA RTX A6000 GPUs with a batch size of 1024. The learning rate schedule follows a cosine decay, starting at 2e-3 and decreasing to a minimum of 2e-4, with the AdamW optimizer (weight decay of 0.05) and cross-entropy loss.
The input features from preprocessing have a size of 2562 with 1 block and 16 output channels, passing through the network composed of 1 Conv Stem Layer and 3 Down Sample AFIDAF Blocks, concluding with the output. The Conv Stem Layer yields an output size of 642, encompassing 4 blocks, and 32 output channels. The first Down Sample AFIDAF Block produces an output of 322 with 2 blocks, and 96 channels for AFIDAF-T (128 for AFIDAF). The second Down Sample AFIDAF Block outputs 162 with 4 blocks and 160 channels for AFIDAF-T (256 for AFIDAF). The third Down Sample AFIDAF Block results in an output size of 82 with 3 blocks, and 192 channels for AFIDAF-T (320 for AFIDAF).
@@ 4.1.1 Results
4.1.2 Ablation on Alternating Domain Filtering
To validate the effectiveness of our alternating Fourier and image domain filtering approach, we compare AFIDAF with AFFNet [12] and IDAF (replacing the AFF block with image domain LKA [9]) on ImageNet-1K. The results, shown in the last 3 lines of Tab. 1, demonstrate the superiority of our alternating domain approach over single-domain methods.
@@ 4.1.2 Ablation on Alternating Domain Filtering
Table 1: Lightweight network classification comparison on ImageNet-1K dataset. IDAF (image domain adaptive filtering only) replaces AFF block’s channel mixing with image domain LKA [9].
@@ 4.2 Object Detection
Table 2: Comparison of AFIDAF variants Object detection on MS-COCO 2017 dataset.
@@
Table 3: AFIDAF variants vs. other LW backbones Semantic segmentation on PASCAL VOC 2012 dataset.
4.2.1 Setting
4.2.2 Results
As shown in Tab. 3, the detection models equipped with AFIDAF outperform other lightweight transformer-based detectors in terms of mean Average Precision (mAP). Specifically, AFIDAF surpasses the second-best AFFNet [12] by 1.8% in mAP. Consistently, AFIDAF-T edges out AFFNet-T by 0.1% in mAP with 0.1M more parameters.
4.3 Semantic Segmentation
4.3.1 Setting
We use the DeepLabv3 [2] framework for semantic segmentation with AFIDAF and AFIDAF-T backbones. Images are resized to $512\times 512$, and models are initialized with ImageNet-1K pretrained weights. Models are trained for 50 epochs on the VOC dataset, using a cosine learning rate scheduler with a base rate of $5\mathrm{e}{-4}$, a minimum rate of $1\mathrm{e}{-6}$, and optimizer AdamW with a weight decay of 0.05. The loss function employed is cross-entropy loss.
4.3.2 Results
In Tab. 3, AFIDAF demonstrates superior performance compared to other lightweight networks for semantic segmentation. Specifically, AFIDAF achieves a mean Intersection over Union (mIoU) of 81.6%, surpassing the second-best lightweight network, AFFNet, by 1.1%. Additionally, AFIDAF-T exceeds AFFNet-T by 1.8% in mIoU.
Model Params (M) Flops (G) Top-1 (%) Top-5 (%)
Swin-T[24] 28 4.5 81.2 95.5
SpectFormer-XS[32] 20 4.0 80.2 94.7
PoolFormer-S12[45] 12 1.8 77.2 -
PoolFormer-S24[45] 21 3.4 80.3 -
GFNet-XS[34] 16 2.9 78.6 94.2
HAFIDAF 14.8 4.45 79.8 95.0
Table 4: Comparison of HAFIDAF with middleweight networks on ImageNet-1K classification. HAFIDAF achieves competitive performance with fewer parameters.
Model Param APbox AP50box AP75box
(M) (%) (%) (%)
R-50[10] 82 46.3 64.3 50.5
DeiT-S[36] 80 48.0 67.2 51.7
Swin-T[24] 86 50.5 69.3 54.9
HAFIDAF 72 48.9 67.6 53.4
Table 5: Object detection performance on COCO dataset. HAFIDAF maintains competitive performance with significantly fewer parameters compared to larger models.
Model Param mIoU mAcc aAcc
(M) (%) (%) (%)
@@ 4.3.2 Results
Table 6: Semantic segmentation performance on Pascal VOC 2012 dataset. HAFIDAF outperforms Swin-T baseline across all metrics with 24% fewer parameters.
4.4 Experimental Evaluation of HAFIDAF
We conduct experiments to evaluate our proposed Hierarchical AFIDAF (HAFIDAF) model, comparing it with state-of-the-art vision transformers models across image classification, semantic segmentation, and object detection tasks.
4.4.1 Image Classification
Table 6 compares HAFIDAF with other middleweight networks on ImageNet-1K. Based on the Swin-T architecture, HAFIDAF reduces parameters by 47% (14.8M vs. 28M) while only decreasing Top-1 and Top-5 accuracy by 1.4% and 0.5%, respectively. This demonstrates HAFIDAF’s efficiency in balancing model size and performance against recent Vision Transformers.
4.4.2 Object Detection
We evaluate HAFIDAF using the Cascade Mask R-CNN framework (Tab. 6). With consistent training settings across models, HAFIDAF achieves a 17% reduction in model size compared to Swin-T, with only a 1.6% drop in APbox. This showcases HAFIDAF’s ability to balance compression and accuracy in detection tasks.
4.4.3 Semantic Segmentation
Using the UperNet framework, we compare HAFIDAF and Swin-T on the Pascal VOC dataset (Tab. 6). HAFIDAF reduces parameters by 24% while improving all three accuracy metrics (mIoU, mAcc, and aAcc), highlighting its effectiveness in dense prediction tasks.